OpenStudy (anonymous):

I need help with functions! Given f(x)=x^2 + 2x, solve for x where f(x+2) = 0

OpenStudy (anonymous):

(x+2)^2 + 2(x+2) (x+2)(x+2) + (2x+4) (x^2+4x+4) + (2x+4) x^2 +6x + 8 = 0 factor. (x+4)(x+2) = 0 x = -4 and -2 f(x+2) = 0 @ -4 and -2

OpenStudy (anonymous):

oh so you just treat f(x+2) like any other number and just plug it in for x?

OpenStudy (anonymous):

yes, just plug (x+2) anywhere you see an "x" in the function :D

OpenStudy (anonymous):

okay haha thanks for the help!